OpenStudy (anonymous):

Find each of the following. f(x) = |x + 3| g(x) = -x^2 (gf)(3) = (fg)(8) = (ff)(1) =

OpenStudy (anonymous):

for the first i got 54 but it's wrong. what am i doing wrong? this is how i did the first: -3^2(3+3)=9(6)=54

OpenStudy (anonymous):

My latest answer is 12 but thats wrong too

OpenStudy (anonymous):

for the first one of course not the whole thing

OpenStudy (anonymous):

Is g(x) = -(\(x^2\)) or (\(-x)^2\)? g(f(3)) = g(|x+3|) = g(|3+3|) = g(6) = -(6)^2 = -36

OpenStudy (anonymous):

it's just -x^2 without any parenthesis.

OpenStudy (anonymous):

i have to add my own parenthesis? i got 36 earlier because i did -6 times -6, which is 36.

OpenStudy (anonymous):

It depends on which way you read it. Using the BEMDAS / PEMDAS / BIRDMAS rules, you're supposed to solve the exponents / indices / roots first, and then do multiplication. So when you have -x^2 you're supposed to do the x^2 part first, and then multiply the whole thing by -1.

OpenStudy (anonymous):

ok ok nice profile pic by the way haha

OpenStudy (anonymous):

Thank you. Psyence rules.

OpenStudy (anonymous):

lolz pokemon's my sh*t haha ...science and math... kind of, sort of. haha